# CodeTaxo: Enhancing Taxonomy Expansion with Limited Examples via Code Language Prompts

### v1

@@ Abstract.
Taxonomies play a crucial role in various applications by providing a structural representation of knowledge. The task of taxonomy expansion involves integrating emerging concepts into existing taxonomies by identifying appropriate parent concepts for these new query concepts. Previous approaches typically relied on self-supervised methods that generate annotation data from existing taxonomies. However, these methods are less effective when the existing taxonomy is small (fewer than 100 entities). In this work, we introduce CodeTaxo, a novel approach that leverages large language models through code language prompts to capture the taxonomic structure. Extensive experiments on five real-world benchmarks from different domains demonstrate that CodeTaxo consistently achieves superior performance across all evaluation metrics, significantly outperforming previous state-of-the-art methods. The code and data are available at https://github.com/QingkaiZeng/CodeTaxo-Pub.
Taxonomy Expansion; Large Language Models; In-context Learning
isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Taxonomy has a hierarchical structure that represents the hypernym-hyponym relations or “is-A” relations between concepts or entities. Researchers have been using relational knowledge to identify semantic relevance for web search (Yin and Shah, 2010; Liu et al., 2020), personalized recommendation (Zhang et al., 2014; Tan et al., 2022), and question answering (Yang et al., 2017). Existing taxonomies are constructed mainly by human experts or through crowd-sourcing, but this manual approach is time-consuming, labor-intensive, and often limited in knowledge coverage (Bordea et al., 2016; Jurgens and Pilehvar, 2016). As new entities continue to emerge, it becomes essential to keep these taxonomies up-to-date by enriching them with these emerging entities. To address the challenges of manual taxonomy construction and maintenance, taxonomy expansion tasks have been developed to automatically expand the existing taxonomies by integrating emerging entities into the relational graph (Jurgens and Pilehvar, 2016).
Recent works primarily contribute to taxonomy expansion following the discriminative method illustrated in Figure 1(a). These methods extract hierarchical information from the existing taxonomy and employ various techniques to model the structural information, such as local Egonet (Shen et al., 2020), mini-paths (Yu et al., 2020) and Ego-Tree (Wang et al., 2021). As pre-trained language models (PLMs) that have enhanced the performance of various text processing systems, PLMs are leveraged to encode entities’ textual descriptions (term and definition) to enrich the semantic representation of entities and improve the performance in taxonomy expansion tasks (Wang et al., 2021, 2022; Liu et al., 2021b; Xu et al., 2022).
@@
(a) Discriminative Methods: A scoring function is trained to evaluate and select the best parent entity from the current taxonomy for the query entity based on the scoring results (Shen et al., 2020; Liu et al., 2021b; Wang et al., 2021).
@@
(b) Generative Methods: Prompting LLMs to generate the parent entity from the current taxonomy based on the query entity.
Figure 1. Two Types of Methods for Taxonomy Expansion
@@
Figure 2. The overview of the framework for CodeTaxo: CodeTaxo reformulates the task of integrating a query entity $q$ into an existing taxonomy $\mathcal{T}_{0}$ as a code completion task using code-based prompts for LLMs.
The primary methodology adopted by many of the aforementioned methods involves learning hierarchical structural information from annotations generated in a self-supervised manner based on existing taxonomies. However, these methods often encounter limitations when the existing taxonomies are small. In such scenarios, the limited size of the taxonomy fails to generate sufficient self-supervised annotation data, which is crucial for training models to effectively learn the hierarchical structure necessary for taxonomy expansion.
@@
Generative Large Language Models (LLMs) such as GPT-4 (Achiam et al., 2023) and Llama family (Touvron et al., 2023; Dubey et al., 2024) have recently demonstrated remarkable capabilities in text comprehension and generation, making them highly effective for tasks aimed at generating structural knowledge (Ye et al., 2022; Bi et al., 2024; Sun et al., 2024b, a). Moreover, increasing the number of parameters in LLMs significantly enhances their generalization capabilities, surpassing smaller pre-trained models, and enabling superior performance in few-shot or zero-shot settings. Consequently, even in scenarios where existing taxonomies cannot provide sufficient self-supervised annotation data for training, LLMs can still effectively perform the task by leveraging the extensive knowledge embedded within their parameters, acquired from large-scale pre-training corpora. In Figure 1(b), we illustrate the pipeline to how generative methods are applied to the taxonomy expansion task.
@@
However, two key challenges arise when employing LLMs for taxonomy expansion tasks. First, unlike traditional NLP tasks such as question answering and machine translation, which naturally lend themselves to text-to-text generation, the taxonomic structure required for this task is often difficult to represent using natural language. In particular, the taxonomic structure to be generated must be serialized into text, which involves “flattening” the taxonomy into a sequence of parent-child entity pairs (Madaan et al., 2022). However, LLMs are primarily pre-trained on unstructured text, making these serialized structured outputs considerably different from the majority of their pre-training data. Moreover, in natural language, semantically related words are usually located within a relatively small range. In contrast, when representing a taxonomy as a linear sequence, entities that are conceptually related may be positioned at greater distances from each other. Second, for large-scale taxonomies, including every entity from the existing taxonomy within the prompt is impractical due to the limited contextual window size of current LLMs and the associated computational costs. Even if we could incorporate all these thousands of entities directly in the prompts, it would result in structural information loss, thereby reducing the clarity with which LLMs can distinguish entity-specific details.
To address these two challenges, in this work, we proposed CodeTaxo, a novel approach to taxonomy expansion that leverages code languages as prompts. Recent progress in large language models of code (Code-LLMs) (Roziere et al., 2023) has demonstrated the potential to employ Code-LLMs for different structure prediction tasks (Madaan et al., 2022; Li et al., 2023; Wang et al., 2023; Li et al., 2024; Bi et al., 2024). The core philosophy of these approaches is that code language can more effectively represent structural data compared to natural language. In CodeTaxo, we reformulate the taxonomy expansion task as a code completion problem to exploit the benefits of code languages. We first define a base class Entity, which encapsulates the surface name and definition of entities to represent their semantic information. Additionally, Entity class includes references to the parent entity and a list of child entities, thereby capturing the taxonomic relations among the entities. Two methods for modifying the taxonomic relations between entities are also included in the Entity class. To represent the existing taxonomy that we aim to expand, we instantiate the Entity class for each entity within the current taxonomy. To address the limitation posed by the contextual window size of LLMs, which cannot accommodate the entire taxonomy, we selectively include only entities that exhibit high similarity to the query entity in the prompt. For encoding the textual descriptions of entities, we utilize SimCSE (Gao et al., 2021), measuring similarity through cosine similarity.
@@
The effectiveness of CodeTaxo has been validated through extensive experiments on two sets of small-scale sub-taxonomies from WordNet (Bansal et al., 2014), Graphine (Liu et al., 2021a), and three large-scale taxonomies from SemEval-2016 (Bordea et al., 2016). The experimental result demonstrates that CodeTaxo in one-shot setting outperforms all self-supervised baseline methods, despite these baselines being well-trained on self-supervised annotation data derived from large-scale taxonomies in SemEval-2016. Specifically, our CodeTaxo in the one-shot setting shows relative improvements in accuracy of 10.26%, 8.89%, and 9.21% on SemEval-Sci, SemEval-Env, and SemEval-Food, respectively. Additionally, we evaluated CodeTaxo using various open-source LLMs, revealing several interesting observations discussed in this work.
In summary, this study makes the following contributions:
@@
  • We develop a similarity-based filter, which employs a small pre-trained model to encode the textual descriptions of entities, ensuring that only highly relevant entities are included in the prompt in relation to the query entity.
@@
Scope and Limitation. This study represents an initial effort to utilize LLMs for taxonomy expansion. Our primary objective is to identify an effective in-context learning strategy to leverage the potential of LLMs. We acknowledge that the performance and scalability of CodeTaxo are constrained by the inherent knowledge of LLMs and the limitations of their context window size. While this paper does not address the challenges of expanding LLM knowledge or increasing context window size, we hope that our work will inspire further research in these areas.
2. Problem Definition
Figure 3. Prompt Overview of CodeTaxo
In this section, we present the key concepts used in this paper and formally define the taxonomy expansion problem:
Definition 2.1 (Taxonomy).
We follow the definition of taxonomy in (Jiang et al., 2023). A taxonomy $\mathcal{T}=(\mathcal{E},\mathcal{H})$ is a tree-like structure, where each entity $e\in\mathcal{E}$ is a conceptual entity, and each edge $h\in\mathcal{H}$ represents the hypernymy-hyponymy relation between the two entities connected by it. Each entity $e$ is associated with a set of textual description $\mathcal{X}_{e}=\{X^{t}_{e},X^{d}_{e}\}$, where $X^{t}_{e}$ is its term and $X^{d}_{e}$ is its definition. Meanwhile, each directed edge $h=\langle p,c\rangle\in\mathcal{H}$ represents a parent-child relationship that points to a child entity $c$ from its most exact hypernymy entity $p$.
Since the taxonomies are naturally incomplete and always changing as emerging entities are introduced. Adding these emerging entities to the existing taxonomy is essential to keep it up-to-date. This study focuses on how to expand taxonomies with given emerging entities. We define the task of taxonomy expansion as follows:
Definition 2.2 (Taxonomy Expansion).
@@ 3.1. Task Instruction
Figure 4. Python code in CodeTaxo defining a Entity class with two methods for managing parent-child relations.
3.2. Taxonomy Code Representation
To represent the existing taxonomy $\mathcal{T}_{0}=(\mathcal{V}_{0},\mathcal{E}_{0})$ as code language, we concatenate the entity class definition, representation of existing taxonomic relations, and the code completion prompt. In this work, we use Python as the programming language for the code prompt due to its widespread popularity.
3.2.1. Entity Class Definition
First, we define a base type Entity to be inherited by each entity mentioned in the taxonomy expansion task. As shown in Figure 4, the code defines a Python class named Entity that models a taxonomic structure with parent-child relations. The first line imports the List type from the typing module, which is used for type hinting. This allows the child attribute to be explicitly declared as a list of Entity objects.
@@ 3.2.1. Entity Class Definition
  • name: A string representing the term of the entity.
  • description: A string representing the textual description of the entity
@@
  • child: A list of entities, each an instance of the Entity class, representing the entity’s children.
@@
3.2.2. Representing the Existing Taxonomy
@@
where $p_{e}$ denotes the parent entity of $e$, and $\mathcal{C}_{e}=[c^{1}_{e},c^{2}_{e},\dots,c^{n}_{e}]$ represents the list of its child entities.
3.2.3. Semantic Similarity Filter
Including every entity $e\in\mathcal{E}_{0}$ in the code prompt to represent the existing taxonomy $\mathcal{T}_{0}$ presents two major challenges. First, the limited contextual window size of LLMs makes it impractical for large-scale taxonomies. Second, it unnecessarily expands the search space, introducing irrelevant entities and redundant information. To mitigate these issues, we propose a Semantic Similarity Filter that selects only entities relevant to the query $q$ for inclusion in the prompt context.
To compute the similarity between a query entity $q$ with its descriptive text $\mathcal{X}q=\{X^{t}_{q},X^{d}_{q}\}$ and an entity $e_{i}\in\mathcal{E}0$ with its descriptive text $\mathcal{X}{e_{i}}=\{X^{t}_{e_{i}},X^{d}_{e_{i}}\}$, we employ the pre-trained language model (PLM) as textual encoder. We concatenate the query entity $q$ and the $i$-th entity $e_{i}$ with special tokens [CLS] and [SEP], then encode the sequence using a pre-trained SimCSE model (Gao et al., 2021). SimCSE converts them into $m$-dimensional representation $\mathbf{q},\mathbf{e}_{i}\in\mathbb{R}^{m}$:
[…]
The semantic relevance is calculated using cosine similarity between $\{\mathbf{e}_{i}\}^{n}_{i=1}$ and $\mathbf{q}$. We select the Top-$k$ entities with high similarity with query entity $q$ from the entity set $\mathcal{E}_{0}$ from existing taxonomy $\mathcal{T}_{0}$ as follow:
where $\mathcal{I}$ is the index set of the selected entities $\mathcal{E}_{sel}=\{e_{i}|i\in\mathcal{I}\}$ that represents the existing taxonomy. In this work, $k$ is set to 50% of the entities in the taxonomy.
3.2.4. Code Completion Prompt.
The code completion prompt involves the instantiation of a query entity $q$ as an instance of the Entity class, as defined in Figure 3.2.1. Since the query entity $q$ lacks information about its parent and child entities, it is instantiated as follows:
[…]
Here, the name and description are initialized with the query’s terms $X^{t}_{q}$ and definition $X^{d}_{q}$, respectively, while the parent is set to None, and the child list is empty.
We propose incorporating an optional feature in the code completion prompt: “then generating a comment to explain why it is the parent of the given query node”. This feature allows the LLM to simultaneously generate both the prediction and its rationale, enhancing the explainability of the output and revealing interesting insights, as discussed in Section 4.6.
3.3. Few-shot Demonstration Construction
4. Experiments
#Concepts #Edges Depth WordNet 20.5 19.5 3.0 Graphine 48.2 48.2 4.6 SemEval-Sci 429.0 451.0 8.0 SemEval-Env 261.0 261.0 6.0 SemEval-Food 1,486.0 1,576.0 8.0
Table 1. Statistics of five taxonomy benchmarks. For WordNet and Graphine, we report the average for taxonomies included in these two benchmarks.
@@ 4. Experiments
TaxoExpan (Shen et al., 2020) 27.8 57.6 11.1 54.8 27.6 54.2 19.8 64.8 24.5 65.9
STEAM (Yu et al., 2020) 36.5 68.2 36.1 69.6 34.2 67.0 23.2 62.4 20.3 63.1
HEF (Wang et al., 2021) 53.6 75.6 55.3 71.4 47.9 73.5 16.4 60.3 25.5 66.5
Musubu (Takeoka et al., 2021) 44.9 76.2 45.3 65.4 42.3 72.4 28.5 64.0 35.4 75.2
TEMP (Liu et al., 2021b) 57.8 85.3 49.2 77.7 47.6 81.0 29.4 65.7 35.9 73.8
BoxTaxo (Jiang et al., 2023) 31.8 64.7 38.1 75.4 31.4 66.8 26.4 63.9 29.2 68.2
TaxoPrompt (Xu et al., 2022) 61.4 85.6 57.4 83.6 53.2 83.1 40.3 71.5 33.9 74.4
TaxoInstruct (Shen et al., 2024) 45.9 76.2 48.8 77.2 34.3 70.2 43.3 71.8 31.8 69.0
@@
Table 2. Performance on taxonomy expansion across two small-scale taxonomies (WordNet and Graphine) and three large-scale taxonomies (SemEval2016: science, environment, food). Bold indicates the highest score; underlined indicates the second-highest. All metrics are in percentages (%).
Our proposed CodeTaxo is evaluated on five real-world benchmarks to address these research questions (RQs):
  • RQ1: How does the performance of the CodeTaxo compare to state-of-the-art baselines in taxonomy expansion task?
  • RQ2: How does CodeTaxo perform across different large language models?
  • RQ3: Which components of CodeTaxo most significantly influence the effectiveness of taxonomy expansion, and how can their hyperparameters be optimized?
4.1. Experimental Settings
4.1.1. Datasets

### v2

@@ Abstract
Taxonomies provide structural representations of knowledge and are crucial in various applications. The task of taxonomy expansion involves integrating emerging entities into existing taxonomies by identifying appropriate parent entities for these new query entities. Previous methods rely on self-supervised techniques that generate annotation data from existing taxonomies but are less effective with small taxonomies (fewer than 100 entities). In this work, we introduce CodeTaxo, a novel approach that leverages large language models through code language prompts to capture the taxonomic structure. Extensive experiments on five real-world benchmarks from different domains demonstrate that CodeTaxo consistently achieves superior performance across all evaluation metrics, significantly outperforming previous state-of-the-art methods. The code and data are available at https://github.com/QingkaiZeng/CodeTaxo-official.
1 Introduction
@@
Taxonomies are hierarchical structures encoding hypernym–hyponym (i.e., “is-A”) relations between concepts or entities. Relational knowledge derived from taxonomies has been widely leveraged to identify semantic relevance for web search Yin and Shah (2010); Liu et al. (2020); Kang et al. (2024), personalized recommendation Zhang et al. (2014); Tan et al. (2022); Huang et al. (2019), and question answering Yang et al. (2017). However, existing taxonomies are mainly constructed by experts or through crowd-sourcing, making the process time-consuming, labor-intensive, and restricted in coverage Bordea et al. (2016); Jurgens and Pilehvar (2016). As new entities emerge, continually enriching taxonomies with these additions becomes vital. To address these challenges, taxonomy expansion aims to integrate new entities into existing taxonomies automatically.
@@
(a) Discriminative Methods: A trained scoring function selects the most appropriate parent entity from the taxonomy for a given query entity.
@@
(b) Generative Methods: LLMs generate the parent entity from the taxonomy based on the query entity.
Figure 1: Two Types Taxonomy Expansion Methods
@@
Figure 2: The overview of the pipeline for CodeTaxo: CodeTaxo reformulates the task of integrating a query entity $q$ into an existing taxonomy $\mathcal{T}_{0}$ as a code completion task using code-based prompts for LLMs.
@@
As shown in Figure 1(a), recent taxonomy expansion methods mainly rely on discriminative methods that model hierarchical structures through techniques like Egonets Shen et al. (2020), mini-paths Yu et al. (2020), and Ego-Trees Wang et al. (2021). Although pre-trained language models (PLMs) enhance these methods by encoding entities’ textual descriptions Wang et al. (2021, 2022); Liu et al. (2021b); Xu et al. (2022), their reliance on limited self-supervised annotations often restricts performance. In contrast, Generative Large Language Models (LLMs) such as GPT-4 Achiam et al. (2023) and Llama family Touvron et al. (2023); Dubey et al. (2024) have recently shown remarkable capabilities in text comprehension and generation, making them highly effective for tasks aimed at generating structural knowledge Ye et al. (2022); Bi et al. (2024); Sun et al. (2024a, b). Increasing LLM parameters boosts generalization, surpassing smaller models and enabling superior few-shot or zero-shot performance. Even with limited annotations, LLMs effectively leverage extensive knowledge embedded within their parameters, acquired from large-scale pre-training corpora. In Figure 1(b), we illustrate the pipeline to how generative methods are applied to the taxonomy expansion task.
@@
Two key challenges arise when applying LLMs to taxonomy expansion. First, unlike traditional text-to-text NLP tasks such as question answering and machine translation, representing the taxonomic structure for this task in natural language is inherently challenging. Specifically, the process requires "flattening" the taxonomy into a sequence of parent-child entity pairs Madaan et al. (2022), effectively serializing a hierarchical structure into linear text. This serialized format is notably different from the unstructured text that LLMs primarily encounter during pre-training. Furthermore, while semantically related words in natural language are usually located near each other, linearizing a taxonomy can separate conceptually related entities by significant distances within the sequence. This disparity adds to the difficulty of aligning LLM outputs with the desired structured representation. Second, scaling to large taxonomies amplifies the problem, as including every entity from the existing taxonomy in the prompt is infeasible. The limited contextual window size of current LLMs and the associated computational overhead imposes strict constraints. Even if it were possible to include thousands of entities within a prompt, the resulting structural information loss would impair the clarity of entity-specific distinctions, reducing the model’s capacity to effectively utilize the taxonomy.
To overcome these challenges, we propose CodeTaxo, a novel taxonomy expansion approach that leverages code language as prompts. Code-based representations have shown promise in structure prediction tasks Madaan et al. (2022); Li et al. (2023); Wang et al. (2023); Li et al. (2024); Bi et al. (2024), as code languages provide a more natural format for structural data. In CodeTaxo, we frame taxonomy expansion as a code completion task. We introduce a base Entity class to store entity surface names, definitions, parent references, and child lists, along with two methods for modifying the taxonomic relations between entities. Each existing taxonomy entity is instantiated as a corresponding Entity object. Due to constraints of contextual window size, we apply a similarity-based filter, using SimCSE Gao et al. (2021) to encode textual description for entities, to include only the most relevant entities in the prompt
@@
We evaluate CodeTaxo through extensive experiments on two sets of small-scale WordNet and Graphine sub-taxonomies Bansal et al. (2014); Liu et al. (2021a), as well as three large-scale SemEval-2016 taxonomies Bordea et al. (2016). Our one-shot CodeTaxo surpasses all self-supervised baselines trained on large-scale SemEval-2016 annotations, achieving relative accuracy improvements of 10.26%, 8.89%, and 9.21% on SemEval-Sci, SemEval-Env, and SemEval-Food, respectively. Additionally, we evaluated CodeTaxo using various open-source LLMs, revealing several interesting observations discussed in this work.
In summary, our main contributions include:
@@
  • We develop a similarity-based filter, which employs a small pre-trained model to encode the textual descriptions of entities, ensuring that only highly relevant entities are included in the prompt concerning the query entity.
@@
2 Problem Definition
Figure 3: Prompt Overview of CodeTaxo
Definition 1 (Taxonomy)
We follow the definition of taxonomy in Jiang et al. (2023). A taxonomy $\mathcal{T}=(\mathcal{E},\mathcal{H})$ is a tree-like structure, where each entity $e\in\mathcal{E}$ is a conceptual entity, and each edge $h\in\mathcal{H}$ represents the hypernymy-hyponymy relation between the two entities connected by it. Each entity $e$ is associated with a set of textual description $\mathcal{X}_{e}=\{X^{t}_{e},X^{d}_{e}\}$, where $X^{t}_{e}$ is its term and $X^{d}_{e}$ is its definition. Meanwhile, each directed edge $h=\langle p,c\rangle\in\mathcal{H}$ represents a parent-child relationship that points to a child entity $c$ from its most exact hypernymy entity $p$.
Definition 2 (Taxonomy Expansion)
@@ 3.1 Task Instruction
Figure 4: Python code in CodeTaxo defining a Entity class for managing parent-child relations.
3.2 Taxonomy Code Representation
To represent the existing taxonomy $\mathcal{T}_{0}=(\mathcal{V}_{0},\mathcal{E}_{0})$ as code language, we concatenate the entity class definition, representation of existing taxonomic relations, and the code completion prompt. We use Python as the programming language for the code prompt due to its widespread popularity.
3.2.1 Entity Class Definition
First, we define a base type Entity to be inherited by each entity mentioned in the taxonomy expansion. In Figure 4, we define a Python class named Entity that models a taxonomic structure with parent-child relations. The first line imports the List type from the typing module, which is used for type hinting. This allows the child attribute to be explicitly declared as a list of Entity objects.
@@ 3.2.1 Entity Class Definition
  • name: A string storing the term of the entity.
  • description: A string storing the textual description of the entity
@@
  • child: A list of entities, each an instance of the Entity class, storing the entity’s children.
@@
3.2.2 Representing the Existing Taxonomy
@@
where $p_{e}$ is the parent entity of $e$, and $\mathcal{C}_{e}=[c^{1}_{e},c^{2}_{e},\dots,c^{n}_{e}]$ is the list of its child entities.
@@
TaxoExpan 27.8 57.6 11.1 54.8 27.6 54.2 19.8 64.8 24.5 65.9
STEAM 36.5 68.2 36.1 69.6 34.2 67.0 23.2 62.4 20.3 63.1
HEF 53.6 75.6 55.3 71.4 47.9 73.5 16.4 60.3 25.5 66.5
Musubu 44.9 76.2 45.3 65.4 42.3 72.4 28.5 64.0 35.4 75.2
TEMP 57.8 85.3 49.2 77.7 47.6 81.0 29.4 65.7 35.9 73.8
BoxTaxo 31.8 64.7 38.1 75.4 31.4 66.8 26.4 63.9 29.2 68.2
TaxoPrompt 61.4 85.6 57.4 83.6 53.2 83.1 40.3 71.5 33.9 74.4
TaxoInstruct 45.9 76.2 48.8 77.2 34.3 70.2 43.3 71.8 31.8 69.0
@@
Table 1: Performance on taxonomy expansion across two small-scale taxonomies (WordNet and Graphine) and three large-scale taxonomies (SemEval2016: science, environment, food). Bold indicates the highest score; underlined indicates the second-highest. All metrics are in percentages (%).
3.2.3 Semantic Similarity Filter
Including all entity $e\in\mathcal{E}_{0}$ to represent the existing taxonomy $\mathcal{T}_{0}$ presents two problems. First, large-scale taxonomies overload the LLM’s limited context window. Second, it unnecessarily expands the search space, introducing irrelevant entities and redundant information. To mitigate these issues, we propose a Semantic Similarity Filter that selects only entities relevant to the query $q$ for inclusion in the prompt context.
To compute the similarity between a query entity $q$ with its descriptive text $\mathcal{X}q=\{X^{t}_{q},X^{d}_{q}\}$ and an entity $e_{i}\in\mathcal{E}0$ with its descriptive text $\mathcal{X}{e_{i}}=\{X^{t}_{e_{i}},X^{d}_{e_{i}}\}$, we employ the pre-trained language model (PLM) as textual encoder. We concatenate the query entity $q$ and the $i$-th entity $e_{i}$ with special tokens [CLS] and [SEP], then encode the sequence using a pre-trained SimCSE model Gao et al. (2021). SimCSE converts them into $m$-dimensional representation $\mathbf{q},\mathbf{e}_{i}\in\mathbb{R}^{m}$:
[…]
The semantic relevance is calculated using cosine similarity between $\{\mathbf{e}_{i}\}^{n}_{i=1}$ and $\mathbf{q}$. We select the Top-$k$ entities most similar to query entity $q$ from the entity set $\mathcal{E}_{0}$ in $\mathcal{T}_{0}$ as follow:
where $\mathcal{I}$ is the index set of the selected entities $\mathcal{E}_{sel}=\{e_{i}|i\in\mathcal{I}\}$ that represents the existing taxonomy. $k$ is set to 50% of the entities in $\mathcal{E}_{0}$.
3.2.4 Code Completion Prompt.
The code completion prompt involves the instantiation of a query entity $q$ as an instance of the Entity class, as defined in Section 3.2.1. Since the query entity $q$ lacks information about its parent and child entities, it is instantiated as follows:
[…]
Here, $X^{t}_{q}$ and $X^{d}_{q}$ define the query’s name and description, while None and [] indicate the absence of parent and child entities.
We propose incorporating an optional feature in the code completion prompt: “then generating a comment to explain why it is the parent of the given query node”. This feature allows the LLM to simultaneously generate both the prediction and its rationale, improving explainability and revealing interesting insights, as discussed in Section 4.5.
3.3 Few-shot Demonstration Construction
4 Experiments
4.1 Experimental Settings
Datasets.